OpenStudy (anonymous):

8(x^52^6)^0 simplify

OpenStudy (anonymous):

a^0=1 (any thing)^0=1 use this

OpenStudy (mathmate):

Assuming x does not equal to zero, (x^52^6)^0 evaluates to 1. So the whole expression is just 8*1=8

OpenStudy (anonymous):

so t is 8^1=8

OpenStudy (anonymous):

yes well done

OpenStudy (anonymous):

thanks